OpenStudy (anonymous):

Evaluate f(x) = -2x raised to the second power -4 for x=3????? HELP!!!

OpenStudy (mertsj):

Whatever does that mean????

OpenStudy (mertsj):

\[f(x)=(-2x)^2\]

OpenStudy (mertsj):

\[f(x)=(-2x)^{-4}\]

OpenStudy (anonymous):

No like... hold on ille draw it idk how to type the numbers like that

OpenStudy (anonymous):

try the equation creator

OpenStudy (mertsj):

\[f(x)=(-2x)^2-4\]

OpenStudy (anonymous):



OpenStudy (mertsj):

\[f(3)=-2(3)^2-4=-2(9)-4=-18-4=-22\]

OpenStudy (anonymous):

oh yeah just found that ille use it next time thanks.... but do you have any idea how to do that ??